the area of a rectangle is 130 square feet and the base is 42in, what is the height?

Answer by jim_thompson5910(35256)

A = (bh)/2

130 = (42h)/2

130*2 = 42h

260 = 42h

260/42 = h

h = 130/21

h = 6.190476 (approximately)